1. Which of the following mixtures exerts the highest total pressure in 5.0-L container?

(a) 0.5 mol H2 + 0.6 mol N2 + 0.4 mol Cl2

(b) 0.7 mol H2 + 0.6 mol O2 + 0.1 mol N2

(c) 0.6 mol H2 + 0.3 mol Ar + 0.7 mol He

2.

Which component (H2, O2 or CO2) exerts the highest partial pressure in a gaseous mixture made of 0.50 mol H2, 0.49 mol O2, and 0.48 mol CO2?

3.

Complete the following conversion table:

680.0 mm Hg

14.7 psi

to kPA, torr and asi
Answer #1

Answer 1 -

We know that,

PV = nRT

P = nRT/V

R, T and V are same for each part So, Pressure is directly proportional to moles.

So, the Part in which total moles are greater will have the greater Total Pressure.

(a) 0.5 mol H2 + 0.6 mol N2 + 0.4 mol Cl2

Total Moles = 0.5 mol + 0.6 mol + 0.4 mol = 1.5 mol

(b) 0.7 mol H2 + 0.6 mol O2 + 0.1 mol N2

Total Moles = 0.7 mol + 0.6 mol + 0.1 mol = 1.4 mol

(c) 0.6 mol H2 + 0.3 mol Ar + 0.7 mol He

Total Moles = 0.6 mol + 0.3 mol + 0.7 mol = 1.6 mol

So, Part C will exert more Total Pressure. [ANSWER]

Answer 2 -

Given,

Moles of H2 = 0.50 mol

Moles of O2 = 0.49 mol

Moles of CO2 = 0.48 mol

According to Dalton's Law,

Partial Pressure = Mole Fraction * Total Pressure

Total Pressure will be same for each, the one which have greater mole fraction will exert more partial pressure.

Now,

Mole Fraction of A = Moles of A / Total Moles

So,

Moles fraction of H2 = 0.50 mol / (0.50 mol + 0.49 mol + 0.48 mol) = 0.340

Moles fraction of O2 = 0.49 mol / (0.50 mol + 0.49 mol + 0.48 mol) = 0.333

Moles fraction of CO2 = 0.48 mol / (0.50 mol + 0.49 mol + 0.48 mol) = 0.326

So, H2 will exert more partial pressure. [ANSWER]

Answer 3 -

Given,

680.0 mm Hg

14.7 psi

to kPa, torr and psi

We know that,

1 mmHg = 0.133322 kPa

So, 680.0 mmHg = 680.0 * 0.133322 kPa = 90.6592 kPa [ANSWER]

1 mmHg = 1 torr

So, 680.0 mmHg = 680.0 * 1 torr = 680.0 torr [ANSWER]

1 mmHg = 0.0193368 psi

So, 680.0 mmHg = 680.0 * 0.0193368 psi = 13.149 psi [ANSWER]

1 psi = 6.89476 kPa

So, 14.7 psi = 14.7 * 6.89476 kPa = 101.353 kPa [ANSWER]

1 psi = 51.7149 torr

So, 14.7 psi = 14.7 * 51.7149 torr = 760.21 torr [ANSWER]

1 psi = 51.7149 mmHg

So, 14.7 psi = 14.7 * 51.7149 mmHg = 760.21 mmHg [ANSWER]